The velocity of waves on a string is 240 m/s. If t
Answer #1

distance between nodes = wavelength/2

wavelength = v/f = 240/175 = 1.3714 m = 137.14 cm

distance between nodes = 137.14/2 = 68.57 cm = 0.69m